# Fatigue-PINN: Physics-Informed Fatigue-Driven Motion Modulation and Synthesis

### v1

@@ Abstract
Fatigue modeling is essential for motion synthesis tasks to model human motions under fatigued conditions and biomechanical engineering applications, such as investigating the variations in movement patterns and posture due to fatigue, defining injury risk mitigation and prevention strategies, formulating fatigue minimization schemes and creating improved ergonomic designs. Nevertheless, employing data-driven methods for synthesizing the impact of fatigue on motion, receives little to no attention in the literature. In this work, we present Fatigue-PINN, a deep learning framework based on Physics-Informed Neural Networks, for modeling fatigued human movements, while providing joint-specific fatigue configurations for adaptation and mitigation of motion artifacts on a joint level, resulting in more realistic animations. To account for muscle fatigue, we simulate the fatigue-induced fluctuations in the maximum exerted joint torques by leveraging a PINN adaptation of the Three-Compartment Controller model to exploit physics-domain knowledge for improving accuracy. This model also introduces parametric motion alignment with respect to joint-specific fatigue, hence avoiding sharp frame transitions. Our results indicate that Fatigue-PINN accurately simulates the effects of externally perceived fatigue on open-type human movements being consistent with findings from real-world experimental fatigue studies. Since fatigue is incorporated in torque space, Fatigue-PINN provides an end-to-end encoder-decoder-like architecture, to ensure transforming joint angles to joint torques and vice-versa, thus, being compatible with motion synthesis frameworks operating on joint angles.
@@
Animation, PINNs, 3CC, Deep Learning, Biomechanics
@@ I Introduction
Studing the effects of fatigue in human motion has been the focus of many biomechanical works [1, 2, 3, 4], where the investigation of possible variations in movement patterns and posture and the inevitable decrease in maximum exerted torques and muscle forces as well as changes in the range of joint angles under fatigue conditions, is conducted. In contrast to typical biomechanics research, such as the previously mentioned studies that do not incorporate modeling of fatigued motions, data-driven approaches for estimating musculoskeletal dynamics (e.g. [5, 6]) offer promising and automated solutions for ergonomically-adjusted motion estimation. This includes fatigue-driven motion modeling, which is rather difficult to achieve without knowing the internal human state. Especially, joint torques are inextricably correlated with muscle activity and can be defined as a function of joint angles [7]. Thus, the determination of joint torques is based on joint angles, velocities, and accelerations, i.e. Inverse Dynamics (ID), and vice versa, i.e. Forward Dynamics (FD). Machine and Deep learning approaches (ML/DL) for estimating physical movement parameters, such as joint contact forces (particularly those of the knee) [8, 9], joint torques [10, 6], and muscle forces [11], function as surrogate models for Inverse and Forward dynamics. These approaches can significantly enhance the simulation of physically-plausible 3D humanoid character movements. In recent years, state-of-the-art approaches such as Physics-Informed Neural Networks (PINNs), which embed physics-based domain knowledge into the training process, have been employed to bridge the gap between kinematics and dynamics estimation techniques [12, 13, 14].
@@
To address this gap in literature, we introduce a deep learning framework based on PINNs, to produce fatigued open-type human movements. As open-type human motion we refer to both non-periodical movements and motions not involving human-object interactions. To account for muscle fatigue, we simulate the fatigue-induced fluctuations modulating joint torques of a 3D humanoid character. This simulation is grounded in a PINN adaptation of the Three-Compartment-Controller (3CC) Model [22, 23, 24], 3CC-$\lambda$, a state machine that describes the transition of all muscle motor units of a human limb from one state (compartment) to another, namely active ($M_{A}$), fatigued ($M_{F}$), or resting ($M_{R}$) states. Given that the majority of motion synthesis frameworks produce sequences of joint angles, we devised an end-to-end encoder-decoder-like architecture, namely Fatigue-PINN, to incorporate fatigue effects within the torque space. In this setup, a surrogate DL model for ID converts joint angles into joint torques, then, the 3CC-$\lambda$ model introduces fatigue to generate fatigued torques, which are afterward processed by a surrogate DL model for FD to perform the reverse transformation. Our semi-dynamics approach is joint-specific, meaning that Fatigue-PINN, is applied for estimating fatigue in each joint, ensuring the generation of realistic fatigue-driven animation without explicitly modeling the complex underlying physics that governs the human body and without the acquisition of fatigued motion capture data for training our models. Our contributions are summarized as follows:
  • Introduction of Fatigue-PINN, an end-to-end automated PINN-based framework to model fatigue in open human movements by exploiting internal human state domain knowledge, bypassing the necessity of acquiring fatigued motion capture data. Our model provides joint-specific fatigue configurations allowing adaptation and mitigation of motion artifacts in a semi-automated manner, resulting in more realistic animations.
  • Using 3CC-$\lambda$, a PINN variation of 3CC, to model fatigue in an automated and physically-consistent manner. PINNs exploit physics-based domain knowledge to constraint the estimation of the output, hence accelerating convergence and accuracy in cases of few training data.
@@
  • The end-to-end encoder-decoder-like architecture of Fatigue-PINN, enables its seamless integration into any animation pipeline that operates on joint angles.
@@ II-A Data-Driven Motion Synthesis
In the motion synthesis area, researchers and developers in the graphics, animation, and gaming industries, dive deeply into both deterministic and probabilistic Deep Learning (DL) motion generation techniques. These data-driven approaches leverage human motion capture and movement history data (e.g., previous frames or motion states like joint angles) to produce or predict the pose and/or joint trajectories of a 3D humanoid character [25]. Especially for deterministic motion synthesis, where the synthesized motion sequence converges to a deterministic pose sequence that regresses towards the mean pose of ground-truth, recurrent architectures like LSTM models [26] for typical motion imitation and RNNs enhanced with phase-functioned networks [27, 28, 29] for scene-aware interactions, are favored. The work in [27], introduces the Neural State Machine (NSM), a framework for goal-driven real-time synthesis of 3D character movements and scene interactions. The NSM consists of a gating network and an RNN motion prediction network, where the first accounts for automatic action state transition based on the global phase of the motion signal and the user input (i.e. desired goal). In the subsequent works of the same authors [28, 29], an architecture similar to NSM is employed and enhanced with: i) a local phase feature that exploits the local motion phases of each skeletal segment to dynamically produce their asynchronous motion in character-character, character-object, and character-scene interactions in [28], and ii) a control scheme in [29] for animation layering to produce sophisticated (e.g. martial arts movements) and novel motions based on reference motions, physics-based simulations, input controls, etc. Convolutional Neural Networks (CNNs) have also been employed for synthesizing human motion sequences [30].
@@ II-B Data-Driven Musculoskeletal Dynamics and Kinematics Estimation
In [5] different feed-forward neural network configurations obtained through hyperparameter space exploration were trained and applied to estimate both joint biomechanical parameters (angles, reactions forces, torques) as well as muscle forces and activations in upper extremities. The authors concluded that more complex neural architectures with optimal parameters obtained through hyperparameter space search, lead to more accurate estimations. Furthermore, RNN models pose as the most common approaches in human biomechanics estimation [8, 6, 10]. In particular, in [8], a Bidirectional Long Short-Term Memory network (BiLSTM), augmented with unsupervised domain adaptation layers, was proposed to perform domain alignment between experimental data from various movements, with the goal to simultaneously generalize across multiple actions and align real-to-synthetic data. LSTM has also been utilized for ankle, knee, and hip joint torque estimation in sit-to-stand trials, alongside other machine learning (ML) and deep learning (DL) techniques such as Linear Regression (LR), Support Vector Machines (SVM), and Convolutional Neural Networks (CNN), as documented in [6], with LSTM indicating the best performance for the task. Similarly, LSTM and Gaussian Process Regression (GPR) were employed in [10] to predict lower extremity joint torques during gait.
@@ II-C Physics-Informed Neural Networks
Physics Informed Neural Networks, first introduced in [46, 47], are standard machine/deep learning models that incorporate the underlying physical laws of their training dataset, expressed as partial differential equations (PDEs), in the learning process (i.e. loss function) to address both forward and inverse problems. Forward problems involve approximating the PDEs governing the training dataset to derive solutions, without prior knowledge of the ground-truth, relying solely on the provided input data and respective boundary conditions. State-of-the-art works in fluid [48, 49] and quantum mechanics/computing [50, 51] utilize PINNs to solve forward problems, without explicitly modeling the underlying physics, unlike finite element and conventional physics-based methods. In inverse problems, PINNs leverage physics-based domain knowledge to penalize the estimation of physical quantities (i.e. limiting the solution space), enhancing robustness and generalization in cases of limited data availability as ground truth, where conventional ML/DL methods are rendered ineffective.
PINNs, despite being the revolution in the area of fluid and quantum mechanics, are also starting to be applied for biomechanics estimation tasks [12, 13, 14, 52, 53, 54, 55, 56]. In [54, 12], PINNs with CNNs as a base are used to predict muscle forces and joint angles based on EMGs, with the work in [54] developing a physics-based domain knowledge transfer technique to develop subject-specific PINN approaches. Furthermore, in [56] a distributed physics-informed DL approach that builds local models acting on subdomains of EMG input data to enhance the efficiency and robustness of muscle forces and joint angle estimation, is presented, while in [53] a PINN-based approach for upper limb joint angle estimation under various loads is explored. In the most recent works [52, 55], involving the estimation of muscle parameters and joint angles [52] as well as joint torques [55] of the upper limbs, novel PINN architectures that incorporate Gated Recurrent Units (GRUs), are implemented. GRUs are a variant of RNNs with fewer training parameters and enhanced computational efficiency compared to LSTM models [57, 58], which significantly improve the accuracy of predictions involving time-dependent inputs and also support the learning of long-term time-dependencies. It is also worth mentioning the study in [13], which predicts muscle forces and identifies muscle-tendon parameters from unlabeled EMGs using both explicit and implicit PINN losses (i.e. describing the implicit relationship between muscle forces predicted by a DL model and those calculated by an embedded musculoskeletal model).
@@ III Methodology
Fatigue-PINN is a PINN-based deep learning architecture to produce fatigued human motion. Our framework consists of three joint-specific modules: i) the ID module to account for the transition from the joint kinematics to the joint torque domain, ii) the Fatigue module, a PINN adaptation of the Three-Compartment-Controller, namely 3CC-$\lambda$, to simulate the effects of fatigue on the synthesized motion and iii) an FD module to transform the fatigued torques into fatigued joint angles, as depicted in Figure 1.
@@
In detail, at each motion frame $t$, the ID module is fed with a vector of $N$ joint angles $\mathbf{q}_{t}^{N}$, which in our case corresponds to 32 joint angles including both the joints of upper (shoulder, elbow, wrist) and lower limbs (hip, knee, ankle), pelvis and lumbar and produces a vector of $N$ joint torques $\hat{\mathbf{\tau}}_{t}^{N}$. These torques are subsequently processed through the 3CC-$\lambda$ model (Fatigue module) to generate fatigued torques ($\acute{\hat{\mathbf{\tau}}}_{t}^{N}$), which are then given as input to the FD module to yield the fatigued motion (i.e. fatigued joint angles $\acute{\hat{\mathbf{q}}}_{t}^{N}$).
Even though our ID/FD BiLSTM model accurately estimates multiple joint torques and angles simultaneously (e.g., achieving approximately ${}86\%$ testing accuracy for upper body joint torques and angles as shown in Table II), we opted for a joint-specific approach (as illustrated in Figure 1). This approach allows the neural network to focus all its learning capacity on a single task, resulting in enhanced accuracy, faster convergence, reduced training times, and lower computational requirements due to its reduced complexity. More specifically, the ID/FD model extracts latent space correlations from all joint angles (torques) to effectively produce predictions solely for one joint. This method allows for handling motion artifacts on a joint level while enabling the setting of joint fatigue profiles as discussed in Section III-B.
@@ III-A Dataset & Data Processing
To train our BiLSTM ID and FD models, we utilized the Adapt Emotional Actions open dataset, from the SIG Center for Computer Graphics Multi Modal Motion Capture Library of the University of Pennsylvania [^1] Available online at: https://fling.seas.upenn.edu/~mocap/cgi-bin/Database, that contains human motion capture data in C3D and BVH formats as well as force plate data (i.e. GRF data) from various activities, including knocking (a bell), box lifting, picking up pillow, pointing to spot, punching, box pushing, throwing, walking and waving. Each motion trial is expressed using a different emotion/behavior, based on the OCEAN model that provides a description of one’s personality over five dimensions: Openness (imagination and creativity), Conscientiousness (organization, carefulness), Extroversion (social aspect of human character), Agreeableness (friendliness, generocity, etc.), and Neuroticism (emotional instability, tendency towards negative emotions), with each factor being bipolar (+/+/-+ / -) and composed of several traits as cited in [59]. These personality traits, represented by the acronym OCEAN, are utilized to label each motion trial, meaning that the corresponding movement is performed to convey a specific emotion or behavior. For example, in the motion trial ”Waving_E$+$_02”, the subject waves in a cheerful, energetic manner (i.e. E$+$ stands for Extroversion, thus, social, active, assertive, dominant, energetic behavior). Therefore, each movement consists of trials that are slightly different from one another, due to the expressed emotions, thereby introducing variability to the the dataset within each motion. This variability enhances the generalization capability of our BiLSTM models across different variations of the same movement.
@@
Raw motion capture and GRF data are rotated to align with the OpenSim [45] coordinate system and converted into TRC and MOT formats (motion files’ formats) using Python scripts. These motion files are then utilized in the musculoskeletal modeling process in OpenSim to derive the ground-truth joint kinematics and corresponding joint torques for training our ID and FD models. The musculoskeletal model [^2] Available online at: https://simtk.org/projects/full_body/ utilized in OpenSim is the one introduced in [60], a Hill-type full-body model with muscle-actuated lower limbs (80 muscle-tendon units) and torque-actuated torso/upper limbs (17 ideal torque actuators) with 37 Degrees of Freedom (DoFs) for describing joint kinematics.
@@
The OpenSim analysis pipeline entails scaling the musculoskeletal model to match the anthropometric measurements of the subject used for capturing the Adapt Emotional Actions dataset. These measurements are obtained from the subject’s neutral pose marker data. Subsequently, the inverse kinematics (IK) procedure is employed to calculate joint angles for both the upper (i.e. arm flexion/adduction/rotation, elbow flexion/pronation and wrist flexion/deviation) and lower limbs (i.e. lumbar extension/bending/rotation, pelvis tilt/list/rotation, hip flexion/adduction/rotation, knee angle, ankle angle and subtalar angle) based on the spatial trajectories of the markers. Static Optimization is then performed to estimate muscle forces, which, along with the joint angles, are used to compute joint torques via ID. As for data pre-processing, the obtained joint kinematics and torques were normalized in the range $[0,1]$ using the min-max normalization method.
@@
The Three-Compartment Controller, first introduced in [22], is a state machine that models the grouping of the muscle motor units of a human limb into three muscle activation states, as well as the transitions between them under static, [22] or dynamic load conditions [23, 24, 61, 62]. As depicted in Figure 2, at the beginning of a motion, motor units from the resting ($M_{R}$) state are transitioned to the active ($M_{A}$) compartment to satisfy the target load, $TL$ (”required force” to fulfill a specific task). If more motor units are activated than the ones required to meet $TL$, some motor units may return to the resting state (from the $M_{A}$ compartment to $M_{R}$). This bidirectional transfer of motor units between $M_{A}$ and $M_{R}$ is governed by a time-variant feedback controller, $C(t)$, which is dependent on TL as shown in Equation (4). During activity, motor units inevitably end up at the fatigued state compartment, $MF$ (i.e. their force start to decay over time), at a fatiguing rate, $F$, and, then, they enter rest period, i.e. moved to $MR$, by a recovery rate of $R$. These transitions between the three states of 3CC in dynamic load conditions are described by the Equations (1), (2) and (3) [23, 24, 62]:
Figure 2: 3CC. Figure reproduced from [24]
@@
$C(t)=\begin{cases}L_{D}*(TL-M_{A})&{\scriptstyle\text{if }M_{A}<TL\text{ and }M_{R}>(TL-M_{A})}\\ L_{D}*M_{R}&{\scriptstyle\text{if }M_{A}<TL\text{ and }M_{R}\leq(TL-M_{A})}\\ L_{R}*(TL-M_{A})&{\scriptstyle\text{if }M_{A}\geq TL}\\ \end{cases}$ (4)
The values of $F$ and $R$ are specific for every joint, as can be found in the literature [24] and depend on the task the muscles perform. A greater ratio between parameters $F$ and $R$ accounts for faster accumulation of fatigue as reported in [23, 24, 62]. Moreover, the parameters $L_{D}$ and $L_{R}$, which contribute to the formulation of $C(t)$ in Equation (4), are constants that define respectively the ratio of muscle force development ($L_{D}$), thereby constraining the rate of muscle activation and the rate of muscle relaxation ($L_{R}$), ensuring the system’s stability is preserved [23, 62].
Each one of the three compartments, $M_{A}$, $M_{F}$, and $M_{R}$ are expressed in %$\%MVC$ – maximum voluntary contractions, and by summing the percentage of muscle motor units in each compartment a total of $100\%$ is obtained. Residual Capacity ($RC$) quantifies the remaining muscle strength or maximum exerted force/joint torque capability as a percentage, accounting for the effects of fatigue, and, thus, is defined as follows [23]:

### v2

@@ Abstract
Fatigue modeling is essential for motion synthesis tasks to model human motions under fatigued conditions and biomechanical engineering applications, such as investigating the variations in movement patterns and posture due to fatigue, defining injury risk mitigation and prevention strategies, formulating fatigue minimization schemes, and creating improved ergonomic designs. Nevertheless, employing data-driven methods for synthesizing the impact of fatigue on motion, receives little to no attention in the literature. In this work, we present Fatigue-PINN, a deep learning framework based on Physics-Informed Neural Networks, for modeling fatigued human movements, while providing joint-specific fatigue configurations for adaptation and mitigation of motion artifacts on a joint level, resulting in more smooth, hence physically-plausible animations. To account for muscle fatigue, we simulate the fatigue-induced fluctuations in the maximum exerted joint torques by leveraging a PINN adaptation of the Three-Compartment Controller model to exploit physics-domain knowledge for improving accuracy. This model also introduces parametric motion alignment with respect to joint-specific fatigue, hence avoiding sharp frame transitions. Our results indicate that Fatigue-PINN accurately simulates the effects of externally perceived fatigue on open-type human movements being consistent with findings from real-world experimental fatigue studies. Since fatigue is incorporated in torque space, Fatigue-PINN provides an end-to-end encoder-decoder-like architecture, to ensure transforming joint angles to joint torques and vice-versa, thus, being compatible with motion synthesis frameworks operating on joint angles.
@@
Animation, Biomechanics, Deep Learning, PINNs, 3CC
[…]
=-21pt
@@ I Introduction
Studing the effects of fatigue in human motion has been the focus of many biomechanical works [1, 2, 3, 4], where the investigation of possible variations in movement patterns and posture and the inevitable decrease in maximum exerted torques and muscle forces as well as changes in the range of joint angles under fatigue conditions, is conducted. In contrast to typical biomechanics research, such as the previously mentioned studies that do not incorporate modeling of fatigued motions, data-driven approaches for estimating musculoskeletal dynamics (e.g. [5, 6]) offer promising and automated solutions for ergonomically-adjusted motion estimation. This includes fatigue-driven motion modeling, which is rather difficult to achieve without knowing the internal human state. Especially, joint torques are inextricably correlated with muscle activity and can be defined as a function of joint angles [7]. Thus, the determination of joint torques is based on joint angles, velocities, and accelerations, i.e. Inverse Dynamics (ID), and vice versa, i.e. Forward Dynamics (FD). Machine and Deep learning approaches (ML/DL) for estimating physical movement parameters, such as joint contact forces (particularly those of the knee) [8, 9], joint torques [10, 6], and muscle forces [11], function as surrogate models for Inverse and Forward dynamics. These approaches can significantly enhance the simulation of physically-plausible 3D humanoid character movements. In recent years, state-of-the-art approaches such as Physics-Informed Neural Networks (PINNs) have been employed to bridge the gap between kinematics and dynamics estimation techniques [12, 13, 14].
@@
To address this gap in the literature, we introduce a deep learning framework based on PINNs to produce fatigued open-type human movements. As open-type human motion, we refer to both non-periodical movements and motions not involving human-object interactions. To account for muscle fatigue, we simulate the fatigue-induced fluctuations modulating joint torques of a 3D humanoid character. This simulation is grounded in a PINN adaptation of the Three-Compartment-Controller (3CC) Model [22, 23, 24], 3CC-$\lambda$, a state machine that describes the transition of all muscle motor units of a human limb from one state (compartment) to another, namely active ($M_{A}$), fatigued ($M_{F}$), or resting ($M_{R}$) states. Given that the majority of motion synthesis frameworks produce sequences of joint angles, we devised an end-to-end encoder-decoder-like architecture, namely Fatigue-PINN, to incorporate fatigue effects within the torque space. In this setup, a surrogate DL model for ID converts joint angles into joint torques, then, the 3CC-$\lambda$ model introduces fatigue to generate fatigued torques, which are afterward processed by a surrogate DL model for FD to perform the reverse transformation. Our semi-dynamics approach is joint-specific, meaning that Fatigue-PINN, is applied for estimating fatigue in each joint, ensuring the generation of smooth fatigue-driven animation without explicitly modeling the complex underlying physics that governs the human body and without the acquisition of fatigued motion capture data for training our models. It is worth noting that fatigued motion capture datasets containing open-type actions or more sophisticated movements, which can be used for training fatigued motion synthesis DL models, are absent from the literature. The latter, along with the lack of data-driven models for the generation of fatigued human motion, serve as the motivation behind this work. Our contributions are summarized as follows:
  • Introduction of Fatigue-PINN, an end-to-end automated PINN-based framework to model fatigue in open human movements by exploiting internal human state domain knowledge, bypassing the necessity of acquiring fatigued motion capture data. Our model provides joint-specific fatigue configurations allowing adaptation and mitigation of motion artifacts in a semi-automated manner, resulting in smoother animations.
  • Using 3CC-$\lambda$, a PINN variation of 3CC, to model fatigue in an automated and physically-consistent manner.
@@
  • The end-to-end encoder-decoder-like architecture of Fatigue-PINN enables its seamless integration into any animation pipeline that operates on joint angles.
A demo of Fatigue-PINN is made available at GitHub [^1] https://github.com/loi10/Fatigue-PINN.
@@ II-A Data-Driven Motion Synthesis
In the motion synthesis area, researchers and developers in the graphics, animation, and gaming industries, dive deeply into both deterministic and probabilistic Deep Learning (DL) motion generation techniques. These data-driven approaches leverage human motion capture and movement history data (e.g., previous frames or motion states like joint angles) to produce or predict the pose and/or joint trajectories of a 3D humanoid character [25]. Especially for deterministic motion synthesis, where the synthesized motion sequence converges to a deterministic pose sequence that regresses towards the mean pose of ground-truth, recurrent architectures like LSTM models [26] for typical motion imitation and RNNs enhanced with phase-functioned networks [27, 28, 29] for scene-aware interactions, are favored. The work in [27] introduces the Neural State Machine (NSM), a framework for goal-driven real-time synthesis of 3D character movements and scene interactions. The NSM consists of a gating network and an RNN motion prediction network, where the first accounts for automatic action state transition based on the global phase of the motion signal and the user input (i.e. desired goal). In the subsequent works of the same authors [28, 29], an architecture similar to NSM is employed and enhanced with: i) a local phase feature that exploits the local motion phases of each skeletal segment to dynamically produce their asynchronous motion in character-character, character-object, and character-scene interactions in [28], and ii) a control scheme in [29] for animation layering to produce sophisticated (e.g. martial arts movements) and novel motions based on reference motions, physics-based simulations, input controls, etc. Convolutional Neural Networks (CNNs) have also been employed for synthesizing human motion sequences [30].
@@ II-B Data-Driven Musculoskeletal Dynamics and Kinematics Estimation
In [5], different feed-forward neural network configurations obtained through hyperparameter space exploration were trained and applied to estimate both joint biomechanical parameters (angles, reaction forces, torques) as well as muscle forces and activations in the upper extremities. The authors concluded that more complex neural architectures with optimal parameters obtained through hyperparameter space search, lead to more accurate estimations. Furthermore, RNN models pose as the most common approaches in human biomechanics estimation [8, 6, 10]. In particular, in [8], a Bidirectional Long Short-Term Memory network (BiLSTM), augmented with unsupervised domain adaptation layers, was proposed to perform domain alignment between experimental data from various movements, with the goal to generalize across multiple actions and align real-to-synthetic data simultaneously. LSTM has also been utilized for ankle, knee, and hip joint torque estimation in sit-to-stand trials, alongside other machine learning (ML) and deep learning (DL) techniques such as Linear Regression (LR), Support Vector Machines (SVM), and CNN, as documented in [6], with LSTM indicating the best performance for the task. Similarly, LSTM and Gaussian Process Regression (GPR) were employed in [10] to predict lower extremity joint torques during gait.
@@ II-C Physics-Informed Neural Networks
Physics Informed Neural Networks, first introduced in [46, 47], are standard machine/deep learning models that incorporate the underlying physical laws of their training dataset, expressed as partial differential equations (PDEs), in the learning process (i.e. loss function) to address both forward and inverse problems. Forward problems involve approximating the PDEs governing the training dataset to derive solutions, without prior knowledge of the ground-truth, relying solely on the provided input data and respective boundary conditions. State-of-the-art works in fluid [48, 49] and quantum mechanics/computing [50, 51] utilize PINNs to solve forward problems, instead of finite element or conventional physics-based methods. In inverse problems, PINNs leverage physics-based domain knowledge to penalize the estimation of physical quantities, enhancing robustness, accuracy, and generalization in cases of limited data availability as ground-truth, where conventional ML/DL methods are rendered ineffective.
PINNs, despite being the revolution in the area of fluid and quantum mechanics, are also starting to be applied for biomechanics estimation tasks [12, 13, 14, 52, 53, 54, 55, 56]. In [54, 12], PINNs with CNNs as a base are used to predict muscle forces and joint angles based on EMGs, with the work in [54] developing a physics-based domain knowledge transfer technique to develop subject-specific PINN approaches. Furthermore, in [56] a distributed physics-informed DL approach is presented that builds local models acting on subdomains of the EMG input data to enhance the efficiency and robustness of muscle forces and joint angle estimation, while in [53] a PINN-based approach is explored for estimating the joint angle of the upper limb under various loads. In the most recent works [52, 55], involving the estimation of muscle parameters and joint angles [52] as well as joint torques [55] of the upper limbs, novel PINN architectures that incorporate Gated Recurrent Units (GRUs), are implemented. GRUs are a variant of RNNs with fewer training parameters and enhanced computational efficiency compared to LSTM models [57, 58], which significantly improve the accuracy of predictions involving time-dependent inputs and also support the learning of long-term time-dependencies. It is also worth mentioning the study in [13], which predicts muscle forces and identifies muscle-tendon parameters from unlabeled EMGs using both explicit and implicit PINN losses (i.e. describing the implicit relationship between muscle forces predicted by a DL model and those calculated by an embedded musculoskeletal model).
@@ III Methodology
Fatigue-PINN is a PINN-based deep learning architecture to produce fatigued human motion. Our framework consists of three joint-specific modules: i) the ID module to account for the transition from the joint kinematics to the joint torque domain, ii) the Fatigue module, a PINN adaptation of the Three-Compartment-Controller, namely 3CC-$\lambda$, to simulate the effects of fatigue on the synthesized motion, and iii) an FD module to transform fatigued torques into fatigued joint angles, as depicted in Fig. 1.
@@
In detail, in each motion frame $t$, the ID module is fed with a vector of $N$ joint angles $\mathbf{q}_{t}^{N}$, which in our case corresponds to 32 joint angles including both the joints of upper (shoulder, elbow, wrist) and lower limbs (hip, knee, ankle), pelvis, and lumbar and produces a vector of $N$ joint torques $\hat{\mathbf{\tau}}_{t}^{N}$. These torques are subsequently processed through the 3CC-$\lambda$ model (Fatigue module) to generate fatigued torques ($\acute{\hat{\mathbf{\tau}}}_{t}^{N}$), which are then given as input to the FD module to yield the fatigued motion (i.e. fatigued joint angles $\acute{\hat{\mathbf{q}}}_{t}^{N}$).
Even though our ID / FD BiLSTM model accurately estimates multiple joint torques and angles simultaneously (e.g., achieving approximately $86\%$ testing accuracy for upper body joint torques and angles as shown in Table V), we opt for a joint-specific approach (as illustrated in Fig. 1). This approach allows the neural network to focus all its learning capacity on a single task, resulting in enhanced accuracy, faster convergence, reduced training times, and lower computational requirements due to its reduced complexity. More specifically, the ID/FD model extracts latent space correlations from all joint angles (torques) to effectively produce predictions solely for one joint. This method allows for handling motion artifacts on a joint level while enabling the setting of joint fatigue profiles as discussed in Section III-B.
Supplementary to this Methodology section, is the accompanying Appendix, where more details concerning the procedure we followed to develop and train our Fatigue-PINN framework, are provided.
@@ III-A Dataset & Data Processing
To train our BiLSTM ID and FD models, we utilized the Adapt Emotional Actions open dataset, from the SIG Center for Computer Graphics Multi-Modal Motion Capture Library of the University of Pennsylvania [^2] Available online at: https://fling.seas.upenn.edu/~mocap/cgi-bin/Database.php, that contains human motion capture data in C3D and BVH formats as well as force plate data (i.e. GRF data) from various activities, including knocking (a bell), box lifting, picking up a pillow, pointing to spot, punching, box pushing, throwing, walking and waving. The dataset is recorded over a subject’s $11$ motion trials for each action, hence containing a total of $99$ motion trials, i.e. $9$ actions with $11$ motion trials each.
Each motion trial is expressed using a different emotion/behavior, based on the OCEAN model that provides a description of one’s personality over five dimensions: Openness (imagination and creativity), Conscientiousness (organization, carefulness), Extroversion (social aspect of human character), Agreeableness (friendliness, generosity, etc.), and Neuroticism (emotional instability, tendency towards negative emotions), with each factor being bipolar (+/+/-) and composed of several traits as cited in [59]. These personality traits, represented by the acronym OCEAN, are utilized to label each motion trial, meaning that the corresponding movement is performed to convey a specific emotion or behavior. For example, in the motion trial ”Waving_E++_02”, the subject waves in a cheerful, energetic manner (i.e. E++ stands for Extroversion, thus, social, active, assertive, dominant, energetic behavior). Therefore, each movement consists of trials that are slightly different from one another, due to the expressed emotions, thereby introducing variability to the dataset within each motion. In particular, the standard deviation of the investigated joint angles for each different motion trial, as well as their standard deviation at every time frame across different emotions ranges from approximately 0.01 to 0.4. This variability is small, yet sufficient to enhance the generalization capability of our BiLSTM models across different variations of the same movement.
@@
Raw motion capture and GRF data are rotated to align with the OpenSim coordinate system and converted into TRC and MOT formats (motion files’ formats) using Python scripts. These motion files are then utilized in the musculoskeletal modeling process in OpenSim to derive the ground-truth joint kinematics and corresponding joint torques for training our ID and FD models. The musculoskeletal model [^3] Available online at: https://simtk.org/projects/full_body/ utilized in OpenSim is the one introduced in [60], a Hill-type full-body model with muscle-actuated lower limbs (80 muscle-tendon units) and torque-actuated torso/upper limbs (17 ideal torque actuators) with 37 Degrees of Freedom (DoFs) to describe joint kinematics.
@@
The OpenSim analysis pipeline entails scaling the musculoskeletal model to match the anthropometric measurements of the subject used for capturing the Adapt Emotional Actions dataset. These measurements are obtained from the subject’s neutral pose marker data. Subsequently, the inverse kinematics (IK) procedure is employed to calculate joint angles for both the upper (i.e. arm flexion/adduction/rotation, elbow flexion/pronation, and wrist flexion/deviation) and lower limbs (i.e. lumbar extension/bending/rotation, pelvis tilt/list/rotation, hip flexion/adduction/rotation, knee angle, ankle angle, and subtalar angle) based on the spatial trajectories of the markers. The computed joint angles, along with the experimentally measured GRF data are used to calculate joint torques via ID.
As for data pre-processing, the obtained joint kinematics and torques were normalized in the range $[0,1]$ using the min-max normalization method. Furthermore, it is worth noting that in the raw dataset, the duration (i.e. number of frames) of each motion trial varies, ranging from approximately $450$ to $600$ frames. Therefore, we resample each motion trial to the length of the longest trial using interpolation, ensuring that all emotional variations of an action have the same number of time steps. After interpolation, each motion trial has a duration of $\sim 600$ frames, resulting in actions with motion trials with a total duration of $\sim 6600$ frames.
@@
The Three-Compartment Controller, first introduced in [22], is a state machine that models the grouping of the muscle motor units of a human limb into three muscle activation states, as well as the transitions between them under static, [22] or dynamic load conditions [23, 24, 61, 62]. As depicted in Fig. 2, at the beginning of a motion, motor units from the resting ($M_{R}$) state are transitioned to the active ($M_{A}$) compartment to satisfy the target load, $TL$ (”required force” to fulfill a specific task). If more motor units are activated than the ones required to meet $TL$, some motor units may return to the resting state (from the $M_{A}$ compartment to $M_{R}$). This bidirectional transfer of motor units between $M_{A}$ and $M_{R}$ is governed by a time-variant feedback controller, $C(t)$, which is dependent on TL as shown in Eq. (4). During activity, motor units inevitably end up in the fatigued state compartment, $MF$ (i.e. their force start to decay over time), at a fatiguing rate, $F$, and, then, they enter rest period, i.e. moved to $MR$, by a recovery rate of $R$. These transitions between the three states of 3CC in dynamic load conditions are described by the Equations (1), (2) and (3) [23, 24, 62]:
Figure 2: 3CC. Figure reproduced from [24].
@@
[…]
The values of $F$ and $R$ are specific for every joint, as can be found in the literature [24] and depend on the task the muscles perform. A greater ratio between parameters $F$ and $R$ accounts for faster accumulation of fatigue as reported in [23, 24, 62]. Moreover, the parameters $L_{D}$ and $L_{R}$, which contribute to the formulation of $C(t)$ in Eq. (4), are constants that define, respectively, the ratio of muscle force development ($L_{D}$), thereby constraining the rate of muscle activation and the rate of muscle relaxation ($L_{R}$), ensuring the system’s stability is preserved [23, 62].
Each one of the three compartments, $M_{A}$, $M_{F}$, and $M_{R}$ is expressed in $\%MVC$ – maximum voluntary contractions, and by summing the percentage of muscle motor units in each compartment, a total of $100\%$ is obtained. Residual Capacity ($RC$) quantifies the remaining muscle strength or maximum exerted force/joint torque capability as a percentage, accounting for the effects of fatigue, and, thus, is defined as follows [23]: